OpenStudy (anonymous):

Which equation could be used to determine the number of stamps, y, that Cassie has in her stamp collection after x months if her stamp collector's starter package included 25 stamps and sends her 10 more stamps each month? I need an answer urgently and explanation I am having hard time with this lessons.

OpenStudy (anonymous):

The choices are A: y=25+10x B: y+25=10x C: y=10+25x D: y+25x=10

OpenStudy (anonymous):

Look at your qusetions..its saying she keeps receiving stamps...meaning ADDing stamps to her collection.

OpenStudy (anonymous):

So it would be A ashley?

OpenStudy (anonymous):

So far it could be A or C. Which seems logical here?

OpenStudy (anonymous):

Stamps = 25+10x Like that.

OpenStudy (anonymous):

Yes,A seems correct. Hope i helped.

OpenStudy (texaschic101):

I agree....It is A.

OpenStudy (texaschic101):

for instance, how many stamps will she have in 5 month ? y = 25 + 10(5) y = 25 + 50 y = 75......in 5 months she will have 75